# Integrating Multimodal Data for Joint Generative Modeling of Complex Dynamics

### v2

@@ Appendix B Supplemental Results
Figure A2: (a) Multimodal integration using MTF on functional magnetic resonance imaging (fMRI)+behavioral data significantly helps to improve DS reconstruction compared to just fMRI data alone (unimodal). Results are for $20$ subjects (black lines; blue line = mean), showing geometrical ($D_{stsp}$, left) and temporal ($D_{H}$, right) disagreement between true and reconstructed systems; $p$-values from paired t-tests as indicated. (b) Example of decoded (color-coding of time series) and true (pale background colors) task stages for one brain area and subject. The trained model was freely iterated forward from the first time step of the test set, and task stages were decoded from the simulated activity. Decoding of class label $\hat{l}\in${Rest, Instruction, CRT, CDRT, CMT} is based on the maximum posterior probability, $\hat{l}_{t}={\arg\max}\ p(l_{kt}|\bm{z}_{t})$, given the latent trajectory $\bm{z}_{t}$. Graph below shows true and predicted BOLD signals. (c) 2d subspaces of freely generated latent activity for a dendPLRNN ($M=30,B=10,K=20,\tau=7$) trained jointly on continuous and categorical data by MTF for one example subject. The color coding corresponds to the task labels predicted according to the maximum posterior probability given the latent state as in (b) at each time step. The latent space appears to be structured according to the different task stages.
Figure A3: Impact of different loss terms in eq. 10 on DS reconstruction quality, indicating that all loss components are crucially important. Comparison of state space agreement $D_{stsp}$ (a) when excluding different terms from the total loss in eq. 10, and (b) for different scaling of the consistency loss ($\mathcal{L}_{\text{con}}$), for the dendPLRNN trained by MTF on data from the chaotic Lorenz system as in 3.1.
@@
Figure A5: Example reconstructions (best out of $10$ trained models each) of a dendPLRNN trained via MTF, compared to highly noisy Gaussian and ordinal training data (orange) produced by the Lorenz system. Reconstruction quality crucially depends on the choice of an optimal forcing interval, similar as observed for STF in (Mikhaeil et al., 2022). Choosing the interval too small (left) or too large (right) leads to significantly worse reconstructions.
@@
Figure A6: Top: DS reconstruction from eight ordinal observations with seven levels each produced by the Lorenz-63 system. Example traces of ordinal time series (center), and example reconstruction by MTF (right) ($M=20,B=10,K=15,\tau=10$), preserving the butterfly-geometry of the Lorenz attractor. Bottom: DS reconstruction of the Rössler attractor from from $75$ ordinal observations with two levels each.
Figure A7: Agreement of true and predicted class label probabilities from a symbolic encoding of the reconstructed chaotic Lorenz-63 dynamics. Bins are color and transparency coded according to their respective relative frequencies. For the reconstructed system, category predictions correspond to those with the maximum posterior probability given a freely generated latent sequence (Eq. 15) sampled from a dendPLRNN trained with MTF ($M=30,B=15,K=30,\tau=10$).
@@
Figure A8: Kernel-density estimates (using the default implementation of the kdeplot function in the seaborn library (Waskom, 2021)) of (a) distribution of maximum Lyapunov exponents, and (b) cumulative distributions of geometrical agreement $D_{stsp}$, for reconstructions produced by the MTF, Multiple Shooting (MS), and sequential multimodal VAE (SVAE), all trained with the same dendPLRNN model ($M=30,\ B=15,\ K=30$) on a symbolic encoding of the Lorenz-63 (eq. A.5). While for the MTF the estimates of Lyapunov exponents are centered around the true value ($\lambda_{max}=0.903$), for SVAE and MS the distributions were much farther off.
@@
Figure A9: (a) Freely generated example trajectories and time series from a dendPLRNN ($M=20,\ B=10,\ K=15,\ \tau=10$) trained with MTF jointly on Gaussian, ordinal, and count data sampled from a Lorenz-63 system. (b) Example power spectra (Gaussian data) and Spearman autocorrelation functions (ordinal and count data). Simulated latent trajectories faithfully capture the geometry of the Lorenz attractor, as well as the temporal structure of the ground truth data when projected back into observation space.
Figure A10: Freely generated time series of a dendPLRNN ($M=30,B=15,K=30,\tau=7$) trained with MTF on continuous fMRI signals and categorical task variables from subjects a) #2, (b) #3, (c) #4, and (d) #7. Shown are time series from $10$ brain areas per subject. The trained dendPLRNN captures the overall temporal structure of the complex activity patterns from very short time series ($288$ time steps of training data for each subject).

### v3

@@ Appendix B Supplemental Results
Figure A2: a: Multimodal integration using MTF on functional magnetic resonance imaging (fMRI)+behavioral data significantly helps to improve DS reconstruction compared to just fMRI data alone (unimodal). Results are for $20$ subjects (black lines; blue line = mean), showing geometrical ($D_{stsp}$, left) and temporal ($D_{H}$, right) disagreement between true and reconstructed systems; $p$-values from paired t-tests as indicated. b: Example of decoded (color-coding of time series) and true (pale background colors) task stages for one brain area and subject. The trained model was freely iterated forward from the first time step of the test set, and task stages were decoded from the simulated activity. Decoding of class label $\hat{l}\in${Rest, Instruction, CRT, CDRT, CMT} is based on the maximum posterior probability, $\hat{l}_{t}={\arg\max}\ p(l_{kt}|\bm{z}_{t})$, given the latent trajectory $\bm{z}_{t}$. The graph below shows true and predicted BOLD signals. c: 2d subspaces of freely generated latent activity for a dendPLRNN ($M=30,B=10,K=20,\tau=7$) trained jointly on continuous and categorical data by MTF for one example subject. The color coding corresponds to the task labels predicted according to the maximum posterior probability given the latent state as in b at each time step. The latent space appears to be structured according to the different task stages.
Figure A3: Freely generated time series from $10$ brain areas per subject from subjects #$\#3$ (left) and #$\#7$. (right). The trained DSR model, only iterated by providing an initial state, captures the overall temporal structure of the complex activity patterns even from very short time series.
Figure A4: Impact of different loss terms in Eq. 11 on DS reconstruction quality, indicating that all loss components are crucially important. a: Comparison of state space agreement $D_{stsp}$ when excluding different terms from the total loss in Eq. 11, and b for different scaling of the consistency loss ($\mathcal{L}_{\text{con}}$), for the dendPLRNN trained by MTF on data from the chaotic Lorenz system as in Sect. 3.1.
@@
Figure A6: Illustration of how dramatic changes in long-term dynamics differently impact the SVAE and MTF loss. a: A dendPLRNN successfully trained on multimodal observations from the Lorenz-63 system (Fig. 2) is altered by setting a parameter of the linear self-connectivity $A_{22}>1$, which results in globally diverging dynamics, yet locally still consistent with the Lorenz-63. b: The global divergence is clearly apparent in the MTF training-time trajectories $\bm{Z}$ generated using STF with interval $\tau=10$ (right), within which the DSR model evolves freely, unlike the SVAE (left). This divergence leads to large increases in the MTF training loss (see MTF loss curve for $A_{22}>1$), and hence is strongly penalized. This effect is essentially not present for the SVAE, where the global divergence induces no considerable effect on the training loss. Hence, the mismatch in global (long-term) dynamics remains unrecognized by the SVAE, explaining why the MTF approach is so superior. In fact, as shown at the bottom, at the minimum of the SVAE loss ($A_{22}\approx 0.966$), only short-term predictions are correctly reproduced (left), while the MTF at its minimum ($A_{22}\approx 0.637$) produces trajectories whose long-term dynamics agree in their temporal structure with those of the original Lorenz-63 (right).
@@
Figure A7: Example reconstructions (best out of $10$ trained models each) of a dendPLRNN trained via MTF, compared to highly noisy Gaussian and ordinal training data (orange) produced by the Lorenz system. Reconstruction quality crucially depends on the choice of an optimal forcing interval, similar as observed for STF in Mikhaeil et al. (2022). Choosing the interval too small (left) or too large (right) leads to significantly worse reconstructions.
@@
Figure A8: Top: DS reconstruction from eight ordinal observations with seven levels each produced by the Lorenz-63 system. Example traces of ordinal time series (center), and example reconstruction by MTF (right) ($M=20,B=10,K=15,\tau=10$), preserving the butterfly geometry of the Lorenz attractor. Bottom: DS reconstruction of the Rössler attractor from from $75$ ordinal observations with two levels each.
@@
Figure A9: a: True and predicted class label probabilities (given the maximum posterior probability for a category at each time step) from a freely generated trajectory of a dendPLRNN ($M=30,B=15,K=30,\tau=10$), trained with MTF on the symbolic representation of the chaotic Lorenz-63 dynamics. b and c: Kernel-density estimates of distributions of maximum Lyapunov exponents (b) and cumulative distributions of $D_{stsp}$ (c), comparing training with MTF, Multiple Shooting (MS), and multimodal SVAE across $30$ trained models, all using the same dendPLRNN model. While for the MTF the estimates of Lyapunov exponents are centered around the true value ($\lambda_{max}=0.903$), for SVAE and MS the distributions were much farther off.
@@
Figure A10: a: Freely generated example trajectories and time series from a dendPLRNN ($M=20,\ B=10,\ K=15,\ \tau=10$) trained with MTF jointly on Gaussian, ordinal, and count data sampled from a Lorenz-63 system. b: Example power spectra (Gaussian data) and Spearman autocorrelation functions (ordinal and count data). Simulated latent trajectories faithfully capture the geometry of the Lorenz attractor, as well as the temporal structure of the ground truth data when projected back into observation space.
@@
Figure A13: Reconstruction with missing variables (partial observations). Mean encoded states $\langle\tilde{\bm{Z}}\rangle$ (left) after training a shPLRNN (Hess et al., 2023) with MTF on just one-dimensional time series from the Lorenz-63 system ($x$ coordinate, bottom). Note that the reconstructed attractor’s geometry closely resembles that of the true Lorenz-63 obtained from an optimal temporal delay embedding of the $x$ variable (right; time lag chosen as the first minimum of the mutual information, see Kantz & Schreiber (2004)). This illustrates that MTF can perform a kind of implicit delay embedding when the underlying system is not fully observed.
Figure A14: Reconstruction of a $10$d chaotic Lorenz-96 system solely from ordinal observations with up to $15$ levels using a shPLRNN (Hess et al. (2023); $M=10,L=100,\tau=10$) as DSR model. On average, each variable only occupied 7 unique categories due to the random initialization of the observation model. a, top: Ground truth ordinal time series sampled from a randomly initialized ordinal observation model $p({\bm{o}}_{t}|\bm{x}_{t})$ from ground truth states $\bm{x}_{t}$ of the Lorenz-96 system, and reconstructed (generated) ordinal observations using the trained decoder model $p({\bm{o}}_{t}|\tilde{\bm{z}}_{t})$. Bottom: Example ground truth and freely generated ordinal time series from 1 channel. b: Ground truth states $\bm{x}_{t}$ (top), states encoded using the mean of the states taken from the trained MTF encoder $p(\tilde{\bm{z}}_{t}|\bm{o}_{t})$ (center), and freely generated latent activity from the trained DSR model ${\bm{z}}_{t}=F_{\theta}(\bm{z}_{t-1})$ (bottom). A linear operator $\bm{B}$ was optimized to align ground truth states $\bm{X}$ of the Lorenz-96 (not seen during training!) and states $\tilde{\bm{Z}}$ encoded from the ordinal data via linear regression ($\bm{B}=\left(\tilde{\bm{Z}}^{T}\tilde{\bm{Z}}\right)^{-1}\tilde{\bm{Z}}^{T}\bm{X}$), and used to project the freely generated activity of the shPLRNN into the observation space of the Lorenz-96 system. c, left: Example of ground truth (orange) and freely generated (blue) activity, bearing the same temporal structure. Right: Aligned ground truth ($\bm{x}_{t}$) and encoded latent states ($\tilde{\bm{z}}_{t}$) as in b for one example unit. Note that the ground truth states $\bm{x}_{t}$ have never been seen by the model during training but are only inferred from the ordinal observations via the MTF encoder $p(\tilde{\bm{z}}_{t}|\bm{o}_{t})$, yet still overlap almost perfectly.